OpenStudy (anonymous):

matt travels two and one-half miles to go the library . this is 3 times the distance that jenny travels. how far does jenny travel? is it 5/6 ??

OpenStudy (anonymous):

jenny = 1/3 (matt) = 1/3 * (5/2) = 5/6

OpenStudy (anonymous):

im right??

OpenStudy (anonymous):

Yes